OpenStudy (anonymous):

What is the equation of the line that passes through the points (−2, 3) and (2, 7)? x − y = −1 x − y = −2 x − y = − 5 x − y = − 6

OpenStudy (anonymous):

@imqwerty

imqwerty (imqwerty):

put the points in the equations given in options the one which satisfies both is the answer :)

OpenStudy (anonymous):

idk how to

imqwerty (imqwerty):

or u can do it like this- equation of a line passing through (x1,y1) and (x2,y2) is given by- \[\frac{ y-y_{1} }{ y_{2}-y_{1}}=\frac{ x-x_{1} }{ x_{2}-x_{1} }\]

OpenStudy (anonymous):

ok

imqwerty (imqwerty):

:)

OpenStudy (anonymous):

y2-y1/x2-x1 to find the slope, once found use y-y1=m(x-x1)

OpenStudy (anonymous):

im stuck

OpenStudy (anonymous):

are you stuck at the beginning?

OpenStudy (anonymous):

yea

imqwerty (imqwerty):

ok so we got the equation lets put (x1,y1) and (x2,y2) \[\frac{ y-3 }{ 7-3 }=\frac{ x- (-2)}{ 2-(-2) }=>\frac{ y-3 }{ 4 }=\frac{ x+2 }{ 4 }\]\[y-3=x+2 =>x-y=-5\]

OpenStudy (anonymous):

ok i see it now , i just needed to see how to put it in form

OpenStudy (anonymous):

so its c?

imqwerty (imqwerty):

yes (:

OpenStudy (anonymous):

ty

imqwerty (imqwerty):

np :D